# Fluctuations in random field Ising models

### v1

@@ 6 Appendix: Proof of technical lemmas
Fix $r\in[2,\infty]$. Recall that the $r\to r$ operator norm can be written as
@@
where $r^{\star}\in[1,2]$ is the Hölder conjugate of $r$ that satisfies $\frac{1}{r}+\frac{1}{r^{\star}}=1$ (e.g. see pg 2 of [26]). Let $S_{r}^{n}:=\{\mathbf{v}\in\mathbb{R}^{n}:\|\mathbf{v}\|_{r}=1\}$ be the $\ell^{r}$ sphere in $\mathbb{R}^{n}$. Also, for any $\varepsilon>0$, let $\mathcal{N}(S_{r}^{n},\varepsilon)$ be the $\varepsilon$-covering number of $S_{r}^{n}$, with distance $d(\mathbf{v},\mathbf{v}^{\prime})=\|\mathbf{v}-\mathbf{v}^{\prime}\|_{r}$. Then, repeating the standard volumetric argument (e.g. see Proposition 4.2.12 in [36]) gives the upper bound
@@
Let $\mathbf{v}_{0}\in S_{r}^{n},~{}\mathbf{w}_{0}\in S_{r^{\star}}^{n}$ be vectors such that attains the supremum, i.e. $\|\mathbf{A}_{n}\|_{r}=\mathbf{v}_{0}^{\top}\mathbf{A}_{n}\mathbf{w}_{0}$. Take $\mathbf{v}\in\mathcal{N}(S_{r}^{n},\varepsilon),~{}\mathbf{w}\in\mathcal{N}(S_{r^{\star}}^{n},\varepsilon)$ such that $\|\mathbf{v}-\mathbf{v}_{0}\|_{r}<\varepsilon,~{}\|\mathbf{w}-\mathbf{w}_{0}\|_{r^{\star}}<\varepsilon$. Then, we can write
@@
Step 2: Concentration. We fix $\mathbf{v}\in\mathcal{N}(S_{r}^{n},\varepsilon),\mathbf{w}\in\mathcal{N}(S_{r^{\star}}^{n},\varepsilon)$, and compute the tail bound for $\operatorname{\mathbb{P}}(\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{w}>\delta)$. For notational simplicity, let us define a $n\times n$ matrix $\mathbf{B}_{n}$ by setting $\mathbf{B}_{n}(i,j)=v_{i}w_{j}I(i\neq j)$. For each $1\leq k\leq N$, let
$T_{k}=T_{k}^{\mathbf{v},\mathbf{w}}:=\sum_{i\neq j}Z_{k,i}Z_{k,j}v_{i}w_{j}=\mathbf{Z}_{k}^{\top}\mathbf{B}_{n}\mathbf{Z}_{k}$
be a quadratic form of $\mathbf{Z}_{k}=(Z_{k,1},\ldots,Z_{k,n})$. Then, we can write
$\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{w}=\frac{1}{N}\sum_{1\leq i\neq j\leq n}\sum_{k=1}^{N}Z_{k,i}Z_{k,j}v_{i}w_{j}=\frac{1}{N}\sum_{k=1}^{N}T_{k}.$
As $\{\mathbf{Z}_{k}\}_{1\leq k\leq N}$ are i.i.d, so are $\{T_{k}\}_{1\leq k\leq N}$. By a moment generating function based version of the Hanson-Wright inequality (e.g. see the last equation in Step 4 in the proof of Theorem 1.1 in [33], where we take $A=\mathbf{B}_{n}$ and $S=\sum_{i\neq j}\mathbf{B}_{n}(i,j)Z_{k,i}Z_{k,j}$), we can write
$\operatorname{\mathbb{E}}\exp[\theta T_{k}]\leq\exp\Big{[}K_{0}\theta^{2}\|\mathbf{B}_{n}\|_{F}^{2}\Big{]},\quad\text{for all}\quad 0<\theta\leq\frac{K_{1}}{\|\mathbf{B}_{n}\|}.$
Here, $K_{0},K_{1}>0$ are finite constants depending only on the sub-Gaussian norm. As $T_{k}$ are i.i.d for each $k=1,\ldots,N$, we have
@@
Noting that
$\|\mathbf{B}_{n}\|\leq\|\mathbf{v}\mathbf{w}^{\top}\|+\max_{i=1}^{N}|v_{i}w_{i}|\leq\|\mathbf{v}\|_{r}\|\mathbf{w}\|_{r^{\star}}+1\leq 2$
(by the triangle and Hölder’s inequality), we have $\theta=\frac{n^{1/r}}{\sqrt{N}}\leq 1\leq\frac{K_{1}}{\|\mathbf{B}_{n}\|}$, by choosing $K_{1}\geq 2$ if necessary.  [^1] When $r=\infty$, we define $n^{1/r}=1$. It is easy to check that the arguments in this proof does not change. Also, using Hölder and $|w_{j}|\leq 1$ gives
$\|\mathbf{B}_{n}\|_{F}^{2}\leq\Big{(}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}v_{i}^{2}\Big{)}\Big{(}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}w_{j}^{2}\Big{)}\leq n^{\frac{r-2}{r}}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}|v_{i}|^{r}\right)^{\frac{2}{r}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}w_{j}^{r^{\star}}\leq n^{1-2/r}.$
By taking $\delta^{\prime}=\Delta\sqrt{N}n^{1-1/r}$ for $\Delta>0$ large enough (to be chosen later) , and simplifying (6.3), we get
@@
$\operatorname{\mathbb{P}}\Big{(}\|\mathbf{A}_{n}\|_{r}>\Delta\frac{n^{1-1/r}}{\sqrt{N}}\Big{)}$ $\leq\sum_{\mathbf{v}\in\mathcal{N}(S_{r}^{n},\varepsilon),~{}\mathbf{w}\in\mathcal{N}(S_{r^{\star}}^{n},\varepsilon)}\operatorname{\mathbb{P}}\left(\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{w}>\Delta\frac{n^{1-1/r}}{\sqrt{N}}\right)$
@@
Without loss of generality, suppose $\{Z_{k,i}\}_{1\leq k\leq N,1\leq i\leq n}$ have sub-Gaussian norms bounded by $K_{0}$. Write $W_{i}:=\frac{\sum_{k=1}^{N}Z_{k,i}^{2}}{N}$. As $\{Z_{k,i}^{2}\}_{1\leq k\leq N,1\leq i\leq n}$ is a collection of sub-exponential with mean 1 and sub-exponential norm $K_{0}^{2}$, Bernstein’s inequality (e.g. Theorem 2.8.3 in [36]) gives $\operatorname{\mathbb{P}}(W_{i}>2)\leq e^{-K_{1}N},$ for some constant $K_{1}>0$ that only depends on $K_{0}$. We refer the reader to Section 2 in [36] for details regarding sub-Gaussian and sub-exponential random variables.
For a fixed $1\leq i\leq n$, conditioning on $(Z_{1,i},\ldots,Z_{N,i})$ we have $\mathbf{A}_{n}(i,j)=\frac{1}{N}\sum_{k=1}^{N}Z_{k,i}Z_{k,j}$ are independent and identically distributed for all $j\neq i$. In particular, each $\mathbf{A}_{n}(i,j)\mid(Z_{1,i},\ldots,Z_{N,i})$ has a sub-Gaussian distribution with mean 00 and sub-Gaussian norm $K_{0}\sqrt{\frac{W_{i}}{N}}$. Hence, for any $\theta>0$, we have
$\operatorname{\mathbb{P}}\Big{(}\sqrt{n}\sum_{j\neq i}\mathbf{A}_{n}(i,j)^{2}>\delta\mid Z_{1,i},\ldots,Z_{N,i}\Big{)}$ $\leq e^{-\theta\delta/\sqrt{n}}\Big{(}\operatorname{\mathbb{E}}\Big{[}e^{\theta\mathbf{A}_{n}(i,j)^{2}}\mid Z_{1,i},\ldots,Z_{N,i}\Big{]}\Big{)}^{n-1}$
$\leq\exp\Big{[}-\frac{\theta\delta}{\sqrt{n}}+\frac{K_{2}^{2}n\theta W_{i}}{N}\Big{]},\quad\forall 0\leq\theta\leq\frac{1}{K_{2}^{2}}\frac{N}{W_{i}},$
for some constant $K_{2}$ depending on $K_{0}$. Here, the first inequality follows from Markov’s inequality, and the second uses the sub-Gaussianity (see [36, Proposition 2.5.2(iii)]. In particular, on the event $W_{i}\leq 2$, we can take $\theta=\frac{N}{2K_{2}^{2}}$ to get
@@
$\operatorname{\mathbb{P}}\Big{(}\sqrt{n}$ $\sum_{j\neq i}\mathbf{A}_{n}(i,j)^{2}>\delta\Big{)}\leq\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\Big{[}\operatorname{\mathbb{P}}\Big{(}\sqrt{n}\sum_{j\neq i}\mathbf{A}_{n}(i,j)^{2}>\delta,~{}W_{i}\leq 2\Big{)}+\operatorname{\mathbb{P}}(W_{i}>2)\Big{]}$
@@
We first show the existence of $(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}$. Suppose not, i.e. there exists a vector $\mathbf{v}\neq\mathbf{0}$, $\lVert\mathbf{v}\rVert=1$ such that $(\mathbf{I}_{n}-\mathbf{C}_{n})\mathbf{v}=\mathbf{v}$. Then, $\mathbf{C}_{n}\mathbf{v}=\mathbf{v}$, and we have a contradiction since $11\leq\lVert\mathbf{C}_{n}\rVert_{2}\leq\lVert\mathbf{C}_{n}\rVert_{q}<$ 1. The second inequality follows from Lemma 5.1 and that $q\geq 2$.
@@
[Acknowledgments] The authors would like to thank Debdeep Pati, Roman Vershynin, and Kaizheng Wang for helpful discussions.

### v2

@@ 6 Appendix: Proof of technical lemmas
For notational simplicity, assume that the inverse-temperature parameter is fixed at $\theta=1$. Fix $r\in[2,\infty]$. Recall that the $r\to r$ operator norm can be written as
@@
where $r^{\star}\in[1,2]$ is the Hölder conjugate of $r$ that satisfies $\frac{1}{r}+\frac{1}{r^{\star}}=1$ (e.g. see pg 2 of [36]). Let $S_{r}^{n}:=\{\mathbf{v}\in\mathbb{R}^{n}:\|\mathbf{v}\|_{r}=1\}$ be the $\ell^{r}$ sphere in $\mathbb{R}^{n}$. Also, for any $\varepsilon>0$, let $\mathcal{N}(S_{r}^{n},\varepsilon)\subset S_{r}^{n}$ be an $\varepsilon$-net of $S_{r}^{n}$, with distance $d(\mathbf{v},\mathbf{v}^{\prime})=\|\mathbf{v}-\mathbf{v}^{\prime}\|_{r}$. Then, repeating the standard volumetric argument (e.g. see Proposition 4.2.12 in [46]) gives the upper bound
@@
Let $\mathbf{v}_{0}\in S_{r}^{n},~\mathbf{w}_{0}\in S_{r^{\star}}^{n}$ be vectors that attains the supremum in (6.1) so that $\|\mathbf{A}_{n}\|_{r}=\mathbf{v}_{0}^{\top}\mathbf{A}_{n}\mathbf{w}_{0}$. Take $\mathbf{v}\in\mathcal{N}(S_{r}^{n},\varepsilon),~\mathbf{w}\in\mathcal{N}(S_{r^{\star}}^{n},\varepsilon)$ such that $\|\mathbf{v}-\mathbf{v}_{0}\|_{r}<\varepsilon,~\|\mathbf{w}-\mathbf{w}_{0}\|_{r^{\star}}<\varepsilon$. Then, we can write
@@
Step 2: Concentration. We fix $\mathbf{v}\in\mathcal{N}(S_{r}^{n},\varepsilon),\mathbf{w}\in\mathcal{N}(S_{r^{\star}}^{n},\varepsilon)$, and compute the tail bound for $\operatorname{\mathbb{P}}(\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{w}>\delta)$. For notational simplicity, let us define a $n\times n$ matrix $\mathbf{B}_{n}$ (depending on $\mathbf{G}_{n}$) by setting $\mathbf{B}_{n}(i,j)={\mathbf{G}_{n}(i,j)}v_{i}w_{j}I(i\neq j)$. In the remainder of the proof, all probabilities will be conditioned on the graph $\mathbf{G}_{n}$. For each $1\leq k\leq N$, let
$T_{k}=T_{k}^{\mathbf{v},\mathbf{w}}:=\sum_{i\neq j}Z_{k,i}Z_{k,j}{\mathbf{G}_{n}(i,j)}v_{i}w_{j}=\mathbf{Z}_{k}^{\top}\mathbf{B}_{n}\mathbf{Z}_{k}$
be a quadratic form of $\mathbf{Z}_{k}=(Z_{k,1},\ldots,Z_{k,n})^{\top}$. Then, we can write
$\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{w}=\frac{1}{N}\sum_{1\leq i\neq j\leq n}\sum_{k=1}^{N}Z_{k,i}Z_{k,j}\mathbf{G}_{n}(i,j)v_{i}w_{j}=\frac{1}{N}\sum_{k=1}^{N}T_{k}.$
As $\{\mathbf{Z}_{k}\}_{1\leq k\leq N}$ are i.i.d., so are $\{T_{k}\}_{1\leq k\leq N}$. By a moment generating function based version of the Hanson-Wright inequality (e.g. see the last equation in Step 4 in the proof of Theorem 1.1 in [43], where we take $A=\mathbf{B}_{n}$ and $S=\sum_{i\neq j}\mathbf{B}_{n}(i,j)Z_{k,i}Z_{k,j}$), we can write
$\operatorname{\mathbb{E}}\exp\Big[\theta T_{k}\Big]\leq\exp\Big[K_{0}\theta^{2}\|\mathbf{B}_{n}\|_{F}^{2}\Big],\quad\text{for all}\quad 0<\theta\leq\frac{K_{1}}{\|\mathbf{B}_{n}\|}.$
Here, $K_{0},K_{1}>0$ are finite constants depending only on the sub-Gaussian norm. As $T_{k}$ are i.i.d. for each $k=1,\ldots,N$, we have
@@
Note that Hölder’s inequality gives
$\|\mathbf{v}\|^{2}=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}v_{i}^{2}\leq\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}v_{i}^{r}\Big)^{2/r}n^{1-2/r}=n^{1-2/r},$
and $|w_{i}|\leq 1$ gives $\|\mathbf{w}\|^{2}=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}w_{i}^{2}\leq\sum_{\mathnormal{i}=1}^{\mathnormal{n}}w_{i}^{r^{\star}}=1.$ [^1] When $r=\infty$, we define $n^{1/r}=1$. It is easy to check that the arguments in this proof does not change. Combining these bounds and using $\mathbf{G}_{n}(i,j)\leq 1$ give
$\|\mathbf{B}_{n}\|^{2}\leq\|\mathbf{B}_{n}\|_{F}^{2}\leq\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}v_{i}^{2}\Big)\Big(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}w_{j}^{2}\Big)\leq n^{1-2/r}.$
Since we assume $N\geq n$, we can take $\theta=\frac{n^{1/r}}{\sqrt{N}}\leq\frac{K_{1}}{\|\mathbf{B}_{n}\|}$, by choosing $K_{1}>1$ if necessary. By taking $\delta^{\prime}=\Delta\sqrt{N}n^{1-1/r}$ for $\Delta>0$ large enough (to be chosen later), and simplifying (6.3), we get
@@
$\operatorname{\mathbb{P}}\Big(\|\mathbf{A}_{n}\|_{r}>2\Delta\frac{n^{1-1/r}}{\sqrt{N}}\Big)$ $\leq\sum_{\mathbf{v}\in\mathcal{N}(S_{r}^{n},\varepsilon),~\mathbf{w}\in\mathcal{N}(S_{r^{\star}}^{n},\varepsilon)}\operatorname{\mathbb{P}}\left(\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{w}>\Delta\frac{n^{1-1/r}}{\sqrt{N}}\right)$
@@
For notational simplicity, assume that the inverse-temperature parameter is fixed at $\theta=1$. Without loss of generality, suppose $\{Z_{k,i}\}_{1\leq k\leq N,1\leq i\leq n}$ have sub-Gaussian norms bounded by $K_{0}$. Write $W_{i}:=\frac{\sum_{k=1}^{N}Z_{k,i}^{2}}{N}$. As $\{Z_{k,i}^{2}\}_{1\leq k\leq N,1\leq i\leq n}$ is a collection of sub-exponential random variables with mean 1 and sub-exponential norm $K_{0}^{2}$, Bernstein’s inequality (e.g. Corollary 2.8.3 in [46]) gives $\operatorname{\mathbb{P}}(W_{i}>2)\leq e^{-K_{1}N},$ for some constant $K_{1}>0$ that only depends on $K_{0}$. We refer the reader to Section 2 in [46] for details regarding sub-Gaussian and sub-exponential random variables.
For a fixed $1\leq i\leq n$, conditioning on $(Z_{1,i},\ldots,Z_{N,i})$ (and possibly $\mathbf{G}_{n}$), we have $\mathbf{A}_{n}(i,j)=\frac{\mathbf{G}_{n}(i,j)}{N}\sum_{k=1}^{N}Z_{k,i}Z_{k,j}$ are independent and identically distributed for all $j\neq i$. In particular, each $\mathbf{A}_{n}(i,j)\mid(Z_{1,i},\ldots,Z_{N,i})$ has a sub-Gaussian distribution with mean 0 and sub-Gaussian norm $\mathbf{G}_{n}(i,j)K_{0}\sqrt{\frac{W_{i}}{N}}$. Hence, for any $\Theta>0$, we have
$\operatorname{\mathbb{P}}\Big(\sqrt{n}\sum_{j\neq i}\mathbf{A}_{n}(i,j)^{2}>\delta\mid$ $Z_{1,i},\ldots,Z_{N,i}\Big)\leq e^{-\Theta\delta/\sqrt{n}}\prod_{j\neq i}\Big(\operatorname{\mathbb{E}}\Big[e^{\Theta\mathbf{A}_{n}(i,j)^{2}}\mid Z_{1,i},\ldots,Z_{N,i}\Big]\Big)$
$\leq\exp\Big[-\frac{\Theta\delta}{\sqrt{n}}+\frac{K_{2}^{2}\Theta W_{i}}{N}\sum_{j\neq i}\mathbf{G}_{n}(i,j)\Big],\quad\forall 0\leq\Theta\leq\frac{1}{K_{2}^{2}}\frac{N}{W_{i}},$
for some constant $K_{2}$ depending on $K_{0}$. Here, the first inequality follows from Markov’s inequality, and the second uses the sub-Gaussianity (see [46, Proposition 2.5.2(iii)]). In particular, on the event $W_{i}\leq 2$, we can take $\Theta=\frac{N}{2K_{2}^{2}}$ and bound $\sum_{j\neq i}\mathbf{G}_{n}(i,j)\leq n$ to get
@@
$\operatorname{\mathbb{P}}\Big(\max_{\mathnormal{i}=1}^{\mathnormal{n}}\sqrt{n}$ $\sum_{j\neq i}\mathbf{A}_{n}(i,j)^{2}>\delta\Big)\leq\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\Big[\operatorname{\mathbb{P}}\Big(\sqrt{n}\sum_{j\neq i}\mathbf{A}_{n}(i,j)^{2}>\delta,~W_{i}\leq 2\Big)+\operatorname{\mathbb{P}}(W_{i}>2)\Big]$
@@
We first show the existence of $(\mathbf{I}_{n}-\mathbf{C}_{n})^{-1}$. Suppose not, i.e. there exists a vector $\mathbf{v}\neq\mathbf{0}$, $\lVert\mathbf{v}\rVert=1$ such that $(\mathbf{I}_{n}-\mathbf{C}_{n})\mathbf{v}=\mathbf{0}$. Then, $\mathbf{C}_{n}\mathbf{v}=\mathbf{v}$, and we have a contradiction since $1\leq\lVert\mathbf{C}_{n}\rVert_{2}\leq\lVert\mathbf{C}_{n}\rVert_{q}<1$. The second inequality follows from Lemma 2.1 and that $q\geq 2$.
@@
[Acknowledgments] The authors would like to thank Matthias Löwe, Debdeep Pati, Roman Vershynin, and Kaizheng Wang for helpful discussions.